OpenStudy (anonymous):

Toby deposited $2,103.24 in a savings account that earns 5.1% simple interest. What will Toby’s account balance be in 10 months?

OpenStudy (anonymous):

all though I can not clearely see whether is it per annum . I=0,051/12=0.00425 per month fv=p(1+in) =2103.24(1+0.00425)(10) =2103.24(1.0425) =2192.6277